Solve for x: (1/8) to the 2x-1 (2x-1 is the exponent) = 8 to the x-1 (x-1 is the exponent.
Found 2 solutions by Nate, Prithwis:
Answer by Nate(3500)

(1/8)^(2x-1) = (8)^(x-1)
(8^(-1))^(2x-1) = (8)^(x-1)
(8)^(-2x+1)=8^(x-1)
-2x + 1 = x - 1
-3x = -2
x = 2/3

Answer by Prithwis(166)

(1/8)^(2x-1) = (8)^(x-1)
(8^(-1))^(2x-1) = (8)^(x-1)
(8)^(-2x+1)=8^(x-1)
-2x + 1 = x - 1
-3x = -2
x = 2/3